Weeland is a small country. Its population total is 500. However there are only 200 wage earners. Each wage earners earns an average of $3,000.00 per year and spends a total of $2,500.00 per year buying goods and services of which $300.00 goes to buying imported goods. Each year the country exports a total of $50,000.00 worth of goods and services. The Government income tax rate is 15%. All government money is spent on education and healthcare. There are only two industries in the country that employ every wage earner. Each industry spends $15,000 each year on new tools and equipment. What is the value of Net Exports? Show your work.
Answer #1

Ans.

The country exports total goods and services = $ 50,000.00

The amount of money spent on imports = 200 wage earners * $ 300.00 for buying imported goods

So , Imports = 200 * 300 = $ 60,000.00

Thus , Net exports = Exports - Imports

Net Exports = 50,000.00 - 60,000.00 = -10,000.00